OpenStudy (lgbasallote):

\[\begin{array} \text{Find} \; \mathcal L \lbrace A(t) \rbrace \; \text{where} \; A(t) &=0 \quad 02\end{array}\]

OpenStudy (lgbasallote):

is it just \[\huge \left[-\frac{te^{-st}}{s} - \frac{e^{-st}}{s^2}\right|_1^2\]

OpenStudy (anonymous):

I dont understand

OpenStudy (lgbasallote):

am i right @UnkleRhaukus ?

OpenStudy (unklerhaukus):

i cant remember , can you show all working

OpenStudy (anonymous):

\[\Huge \int\limits_{0}^{1}e^{-st}(0)+\int\limits_{1}^{2}te^{-st}+\int\limits_{2}^{\infty}e^{-st}(0)\]

OpenStudy (anonymous):

it was a paiecewise function .the first and the third integrals are zero. your work looks fine.

OpenStudy (lgbasallote):

^yes that's what i did

OpenStudy (lgbasallote):

so is it right?

OpenStudy (anonymous):

yes it is. just apply the limits.

OpenStudy (lgbasallote):

ahh thank you thank you!

OpenStudy (anonymous):

yw:)

OpenStudy (anonymous):

is that laplace transformation ?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

i got \[\huge -\frac{2e^{-2s}}{s} - \frac{e^{-2s}}{s^2} + \frac{2e^{-s}}{s}\] is that right?

OpenStudy (anonymous):

oh , I studied that yesterday in class and didn't understand what my teacher has wrote ~.~ I really think I must retake calculus 1

OpenStudy (lgbasallote):

i dont get it either hah lol

OpenStudy (unklerhaukus):

the laplace is in the frequency domain, where as the original function is in the time domain

OpenStudy (lgbasallote):

....so am i right? o.O

OpenStudy (anonymous):

@lgbasallote its ok .just if you want to take LCM s^2 and simplify it. and do not worry you will get it i am here :P

OpenStudy (lgbasallote):

but i am right so far right?

OpenStudy (anonymous):

right

OpenStudy (lgbasallote):

thanks!!

OpenStudy (anonymous):

i just wanted to make sure that you applied limits correctly i am getting . \[\Large \frac{e^{-s}}{s}-\frac{2e^{-2s}}{s}+\frac{e^{-s}}{s^2}-\frac{e^{-2s}}{s^2}\]

OpenStudy (anonymous):

@lgbasallote i hope you can simplify it more.

OpenStudy (lgbasallote):

ahh yes i made a careless mistake thanks